A constant force acting on a body of mass 10kg changes its’velocity from70m/s to120 m/s while the body covers a distance of 200 m. Find the acceleration produced by the applied force and find the applied force.
1
Expert's answer
2019-03-18T12:42:39-0400

a) Let's first find the acceleration of the body from the kinematic equation:


"v^2 = v_0^2 + 2as,""a = \dfrac{v^2 - v_0^2}{2s} = \dfrac{(120 \dfrac{m}{s})^2 - (70 \dfrac{m}{s})^2}{2 \cdot 200 m} = 23.75 \dfrac{m}{s^2}."

b) Finally, from the Newton's Second Law of Motion, we can find the applied force:


"F = ma = 10 kg \cdot 23.75 \dfrac{m}{s^2} = 237.5 N."

Answer:

a) "a = 23.75 \dfrac{m}{s^2}."

b) "F = 237.5 N."